Circular No. 7823

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002au IN UGC 5100 Independent discovery reports of an apparent supernova in UGC 5100

have been received from T. Puckett and A. Langoussis, Mountain Town, GA, on images taken with the Puckett Observatory 0.60-m automated supernova patrol telescope; and from W. D. Li, University of California at Berkeley, on behalf of LOTOSS (cf. IAUC 7514), on images taken the Katzman Automatic Imaging Telescope (KAIT). Puckett provides the following position for SN 2002au: α = 9^{h}34^{m}37^{s}.60, δ = +5^{o}50′15′′.7 (equinox 2000.0), which is 16′′.0 west and 13′′.6 south of the center of UGC 5100. Li provides position end figures 37^{s}.54, 15′′.7. Unfiltered CCD magnitude estimates, except photographic magnitudes for the Palomar Sky Survey (PSS): 1951 Jan. 4 UT, [20.0: (PSS, via Puckett); 1992 Dec. 21 and 1998 Dec. 12, [21.0: (PSS, via Puckett); 2000 Mar. 9, [20.5: (Puckett); 2002 Jan. 18.4, [19.0: (Li); Feb. 1.3, hint of presence (Li); 8.3, 17.7: (Li); 9.22 and 10.14, 18.0 (Puckett); 10.3, 17.3: (Li).

SUPERNOVA 2002av IN ESO 489-G007 M. Schwartz, Cottage Grove, OR; and W. D. Li, University of Cali-

fornia at Berkeley, on behalf of LOTOSS, report the discovery of another apparent supernova (mag ∼ 17.7) on unfiltered images taken with the 0.5-m Tenagra III automated telescope on Feb. 8.2 and 9.2 UT. The new object is located at α = 6^{h}08^{m}06^{s}.34, δ = −23^{o}59′13′′.2 (equinox 2000.0), which is 10′′.8 east and 17′′.7 north of the nucleus of ESO 489-G007. A KAIT image taken on Feb. 1.2 already showed a hint of the new object, while an image taken on Jan. 23.2 showed nothing at this position (limiting mag ∼ 18.5).

SUPERNOVA 2002as IN UGC 3418 S. Jha, T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002as (cf. IAUC 7819), obtained by L. Macri on Feb. 9.13 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova near maximum light. The spectrum exhibits a blue continuum with broad P-Cyg Balmer lines. Adopting the NASA/IPAC Extragalactic Database recession velocity for UGC 3418 (6743 km/s) yields an expansion velocity of 8700 km/s for the supernova, measured to the minimum of the Hβ profile.

2002 February 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green